# Low-energy threshold demonstration for dark matter searches in TREX-DM with an 37Ar source produced at CNA HiSPANoS

### v1

@@ 4 Conclusions
The combined GEM-Micromegas readout system provided preamplification factors of , enabling detection of both characteristic 37Ar emissions. Most notably, the successful observation of the 270 eV L-shell peak represents a significant achievement, as this ultra-low-energy signature is typically challenging to detect in gas-based detectors. The spatial uniformity analysis confirms homogeneous detector response across the active volume, providing a sensitivity map. The systematic threshold analysis reveals remarkable detector sensitivity, with energy thresholds extending down to approximately 20 eV at optimal voltage settings. This performance approaches the fundamental limit imposed by single-electron ionization. Such thresholds have a great impact on the sensitivity projections of TREX-DM to light WIMPs, as discussed in detail in [13], potentially providing access to unexplored regions of the parameter space. Future work will focus on using this calibration protocol for background characterization studies at low energies. Additionally, efforts are underway to develop more refined calibration techniques to probe this eV range. A promising approach that has been successfully implemented in the gaseous TPC field [26] is the use of a UV laser to calibrate in the individual electron regime. This would also allow us to accurately study the detection efficiency of our combined GEM-MM detectors to ultra-low-energy events near the energy threshold region, something we could not achieve in this work due to the imprecise determination of the initial 37Ar activity.

### v2

@@ 4 Conclusions
The combined GEM-Micromegas readout system provided preamplification factors of , enabling detection of both characteristic 37Ar emissions. Most notably, the successful observation of the 270 eV L-shell peak represents a significant achievement, as this ultra-low-energy signature is typically challenging to detect in gas-based detectors. The spatial uniformity analysis confirms homogeneous detector response across the active volume, providing a sensitivity map. The systematic threshold analysis reveals remarkable detector sensitivity, with energy thresholds extending down to approximately 20 eV at optimal voltage settings. This performance approaches the fundamental limit imposed by single-electron ionization. Such thresholds have a great impact on the sensitivity projections of TREX-DM to light WIMPs, as discussed in detail in [GEM_2025], potentially providing access to unexplored regions of the parameter space. Future work will focus on using this calibration protocol for precise background characterization studies at low energies. Additionally, efforts are underway to develop more refined calibration techniques to probe this eV range. A promising approach that has been successfully implemented in the gaseous TPC field [NEWS-G_2019] is the use of a UV laser to calibrate in the individual electron regime. This would also allow us to accurately study the detection efficiency of our combined GEM-MM detectors to ultra-low-energy events near the energy threshold region.